Find f(a), f(a + h), and the difference quotient f(a + h) − f(a)/h. where
h ≠ 0.
f(x) = x2 + 5
f(a)=
f(a+h)=
f(a+h)-f(a)/h

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Is there some reason you can't do that?

Answer by stanbon(75887)

Find f(a), f(a + h), and the difference quotient f(a + h) − f(a)/h. where
h ≠ 0.
f(x) = x^2 + 5
f(a)= a^2+5
------------------
f(a+h)= (a+h)^2 + 5 = a^2 + 2ah + h^2 + 5
---------------------------------------
[f(a+h)-f(a)]/h = [2ah+h^2]/h
-----
= 2a + h
--------------
Cheers,
Stan H.
------------